# Jules Drean MIT CSAIL Miguel Gomez-Garcia MIT CSAIL Fisher Jepsen MIT CSAIL Thomas Bourgeat EPFL  and  Srinivas Devadas MIT CSAIL

### v3

@@ 4.1. Hardware Isolation
Memory Isolation Sanctum (costan2016sanctum, ) already presents hardware mechanisms for memory isolation in the context of an untrusted OS. Like RISC-V Physical Memory Protection (PMP) (waterman2015risc, ), Sanctum introduces a mechanism to partition physical memory and prevent a malicious OS from accessing another security domain’s private memory regions, or different enclaves from accessing each other’s private memory. At its core, the mechanism uses two 64-bit registers per core (one for trusted memory and one for untrusted/shared memory) as memory bitmaps. Each bit represents a fixed-size physical memory region (typically 32MB) and functions as an access-rights flag for the currently-running program. Citadel borrows this mechanism for its simplicity and enforces these checks on every memory access, including speculative accesses and for accesses performed during page walks.
Microarchitectural Isolation Following MI6 (bourgeat2019mi6, ) methodology we partition all spatially shared microarchitectural state (LLC, MSHRs…) and flush temporally shared state on context switch (all cores’ pipeline state, L1 caches, TLBs…). We are not vulnerable the DRAM controller side channels as we inherit MI6’s constant time DRAM padding mechanism. Finally, our processor does not support hyper-threading and we are not vulnerable to related side channels such as port contention.
@@ 4.2. Enabling Shared Memory
Figure 2. The Dual Page Table Mechanism
Sanctum also uses a dual-page-table mechanism (see Figure 2) to differentiate between trusted and untrusted memory based on virtual addresses, which we also borrow in Citadel. Two new registers (equivalent to base and bounds) are used to define the trusted virtual memory range (i.e. the enclave’s virtual memory). The SM enforces the invariant that the physical memory mapped as trusted is owned by the corresponding enclave. When a memory access to a virtual address is performed by a program, one of the two page tables needs to be picked to translate the virtual address. If the virtual address falls within the trusted range, the translation is performed using the enclave page table. Another important invariant enforced by the SM is that the enclave page table is always stored within enclave private memory. If the virtual address falls outside the trusted range, it corresponds to an untrusted or shared memory access and the translation is performed using the OS page table. The OS page table is untrusted and stored in OS memory. Note that for every physical address used during address translation, additional bound checks are performed to ensure they belong to the same security domain (enclave or OS memory) as the translated address. This mechanism is especially handy to eliminate speculative transmitters, as there is no need to perform any speculative address translation (which could leak information) to know whether a memory access is trusted or not.
4.3. Restricting Speculative Access to Shared Memory
@@
Figure 3. Changes to Memory Execution Pipeline. New elements are shown in blue.
@@
Modifying the Memory Execution Pipeline Upon entry into the MemExePipeline, checks are performed on memory instructions to decide if the memory operation is safe to translate and then issue, or not. It is important to perform these checks very early as the next steps in the MemExePipeline touch shared microarchitecture (through address translation and cache hierarchy for instance) potentially transmitting information to an attacker. A memory access is deemed safe if it accesses trusted memory OR if it is non-speculative. Because the enclave trusted memory range is defined in virtual memory, this check can be performed as soon as operands (i.e. virtual address) are available. If an instruction tries to access untrusted or shared memory, we check that it is non-speculative by conservatively checking if it has reached the head of the ROB. If the instruction is safe, it is dispatched to the next stage of the MemExePipeline. If it is not safe, the instruction is squashed and a signal is sent to the MemRS indicating that it should re-dispatch it once it reaches the head of the ROB.
@@
Modifying the Memory Reservation Station Our mechanism requires MemRS being able to re-dispatch some instructions once they become non-speculative: untrusted memory loads will indeed typically need to get dispatched twice.
@@
Disabling the Mechanism for the OS Blocking speculation on memory access can have a significant performance impact. The SM makes sure that nonenclave programs (typically the OS and standard usermode programs) are still allowed to speculate for their memory accesses by disabling the trusted memory mechanism.
4.4. Speculative Transmitters in the L1 D Cache
The core-private part of the memory system, including the L1 D cache, is now shared between trusted memory accesses and untrusted or shared memory accesses. While shared memory cannot be accessed speculatively, trusted memory can. It means that shared memory can be speculatively evicted by a speculative trusted memory access. Furthermore, this eviction will be observable to an attacker. The cache coherency protocol is such that writing to a memory location that is cached by another core will take more time than writing to a memory location that is not cached by another core. As a result, a shared L1 D can be leveraged by an attacker to create a speculative transmitter and mount a spectre-style attack. To our knowledge, we are the first to identify this subtle attack vector. Note that this doesn’t only apply to shared cache sets, but it also applies to contention over other shared resources in the L1 D cache (including the pipeline to access the SRAM). Partitioning the L1 D cache would be too costly, instead we bypass the L1 D cache for untrusted or shared accesses and directly send the requests to the LLC.
4.5. Tagged TLB for Correct Address Translations
When using the dual page table mechanism, the core-private part of the memory system is shared between trusted memory accesses and untrusted memory accesses. For functional correctness, we need to make sure that addresses are translated correctly for each domain. This small microarchitectural structure accelerates address translation by caching intermediate results used by the page walker. Unfortunately, only a few upper bits of the virtual address are used at the beginning of a page walk, easily resulting in aliasing across the two domains. We tag the Address Translation Cache by adding a bit to each entry indicating if the cached information applies to the trusted or untrusted memory address translations. Aliasing can also occur in the TLB if the trusted virtual memory range is smaller than the granularity of a virtual memory page (which can be if we use Mega or Giga Pages). We also tag the TLB entries to avoid address mistranslations.
4.6. Miscellaneous
We statically partition MSHRs between two cores and place a fair round-robin arbiter at the entry of the LLC. Because these two mechanisms only marginally affect the latency of the LLC for a 2-core design, we observe negligible ($\approx$ 0.1%) performance impact.
We also expose controls for speculation and predictors to the software through Control Status Registers (CSRs). Specifically, a program is able to deactivate speculation all together (every instruction waits to reach the head of the ROB), delay the issuing of shared memory accesses to the head of the ROB, deactivate the training of predictors (branch predictors and return address stack) and/or their use (and then it falls back to the predictor $pc+4$).
4.7. Building the Software Infrastructure
We are interested in evaluating our platform on a Linux-ready system and for programs that one would be interested in putting in an enclave (e.g. cryptographic libraries). This means a lot of software needs to be developed to be able to launch enclaves from a Linux user mode process.
First we implement the SM at the most privilege level of RISC-V using a mix of assembler and C and add the necessary pieces of a micro-kernel required to boot Linux on top of it. We also implement what we call a Mini-SM which is a downsized copy of the SM private to each enclave. This avoids contention over most SM resources between enclaves and untrusted programs. To protect the SM against transient execution attacks, we disable speculation during SM execution. Second, we develop a Linux Kernel module that allows user programs to request SM calls and set up enclaves, keeping Linux responsible for the resource management. Finally, we implement a Secure Bootloader and an end-to-end attestation mechanism that makes it possible to uniquely identify an enclave, establish a shared secret with it and exchange encrypted private data. Our attestation mechanism makes it possible for an enclave to be placed anywhere in physical memory and still have consistent identity through the measurement of the direct mapping between virtual addresses and memory content without measuring physical addresses. The unique mapping is attested and enforced by the SM during enclave setup. More implementation details can be found in the evaluation section (see section 6).
5. Citadel 2.0: Improving Performance
Our first baseline successfully achieves security. In this section, we show several mechanisms to improve its performance. First. fixed LLC partitioning á la MI6 destroys performance for our 2-core machine. We show a reconfigurable LLC partitioning scheme and a fine grain flushing mechanism. Furthermore, disabling speculative shared memory accesses makes software requirements especially light, but provides limited performances. We show a relaxed speculative model where we enable pipelined accesses to shared memory but disable predictors and hard-code them to $pc+4$. When this mode is enabled, the program needs to verify more requirements, i.e., that no secret is leaked when speculating $pc+4$. We show that this new mode improves performance for code segments with burst shared memory accesses, but it hampers performances in the general case. We observe that this is due to the TLB being trashed by misspeculations across functions context. This is an effect that was studied (in a nonsecurity context) in the early 2000s(onur1, ; onur2, ; onur3, ): misspeculations can induce large overhead not only due to the delay in executing the right path, large delays are induced due to the trashing of the TLBs. Finally, we show how insights from building the SM make it possible to securely optimize context switch and enable frequent interrupts.
5.1. Evaluation-Guided Optimisations
Our secure baseline runs on FPGA with a Linux environment (see section 6 for our full configuration) which makes it possible to analyses performance bottlenecks on real programs. For this section, we use an elliptic curves (Ed25519) cryptographic library as a running example (see section 6 for more details). At a high level, our enclave interactively receives messages to sign from a benchtest program using shared memory. Out-of-the box, we see an underwhelming overhead of 120% overhead compared to directly (and insecurely) linking the cryptographic library to our benchtest program.
5.2. Reconfigurable LLC Partitioning
An obvious bottleneck in our design is cache partitioning. Sanctum and MI6’s mechanism (i.e. dividing the LLC in 64 equal size slices to match the 64 DRAM regions) is a terrible idea for our system. Our 2-core prototype has a 1MB LLC (see 5), which divided by 64 gives us 16KB LLC slices. This is smaller than the 64KB L1 D+I caches. Because our caches are inclusive, this means an enclave or a program that resides in one memory region cannot even fully use the L1 caches. MI6 circumvented that issue by assuming a 16-core design with a large 16MB LLC.
We build a mechanism for the OS to reconfigure the LLC partitions to improve enclave and other program’s performance. This costly operation is only required when changing the LLC layout, which only happens before launching an application or an enclave with an important memory footprint. The OS requests a reconfiguration through an SM call while no enclaves are running. The SM will first check that the new partitioning is valid, doesn’t affect regions assigned to the SM. It then waits for other cores to flush their L1 caches and to enter a blocked state in the SM, making sure no requests will be sent to the modified regions during reconfiguration (which could put the cache in an incoherent state). The SM will then interact with the hardware to flush the LLC slices that are modified and set the LLC in its new configuration.
At the hardware level, we now design a practical reconfigurable LLC partitioning scheme. We cannot use way partitioning (kiriansky2018dawg, ) as the LLC has only 16 ways, and we need up to 64 slices, and Ironhide’s (ironhide, ) strategy assumes limited reconfiguration. We propose a new, simple, configurable set partitioning scheme that only requires minimal changes in the hardware.
The configuration for each LLC slice (base and bound of the set-index range) is stored in a big register (1216 bits) called the set-index range table. The SM can modify the entries of the set-index range table by sending MMIO requests to the cache, which is seen as a new device. When the LLC receives a request, it will compute the new set index for the corresponding address. We represent the transformation used in Figure 4. First, we extract a region ID from the address upper bits and obtain the corresponding index range base and size by looking up the set-index range table 1. We then extract the address original set index, located in the middle of the address, and tweak it. We compute the original set-index value modulo the size of the index range 2 and add the base index 3. This effectively restricts the new set index to the corresponding index range, partitioning the LLC 4. Note that we must extend the length of the tags used by the LLC to include the entire original index in case an index range is of size 1 and all addresses from the same region map to the same set.
We observe on our cryptographic library that we reduce overheads from 120% to 19.4% by using a non-uniform partition for the LLC. More details can be found in Section 6.5.
Figure 4. LLC set indexing scheme. The function takes an address as an input and output a set index in the LLC slice allocated to its memory region.
LLC fine-grain flush To go hand-in-hand with our reconfigurable LLC partitioning, we also want to be able to selectively flush parts of the LLC. We create a Zero Device in the upper physical address space. It is implemented as a simple piece of logic connected to the cache hierarchy at the same level as the DRAM, but that always sends back zeros. The trick is that the address space covered by the zero device has the same size as the DRAM and is located at addresses that will collide with DRAM addresses in the cache. Once the SM has placed the machine in a correct waiting state (where all the cores have flushed their L1 and are waiting in the SM), it only needs to access an eviction set in the zero-device address space to flush the cache slice. This means that flushing a small LLC slice is much faster than flushing a big one. We look at the performance of our flushing mechanism in the evaluation section (see 6.5).
5.3. Pipelined Shared Accesses
@@
At this step, we observe that our test benchmark still present a 19.4% overhead compared to our insecure benchmark. We decide to measure the performance of different sections of the code to find and analyze the bottleneck. We identify the function creating the slowdown to be located inside the SHA256 library. This function performs burst accesses to shared memory to load the message to be signed. This behavior is exacerbated by the compiler, trying to optimize the code by placing all the load instructions to shared memory one after the other. Indeed, with our current mechanism, memory accesses to shared regions, first cannot use the L1 and have a 10 cycle latency, second cannot be pipelined as each instruction needs to wait to reach the head of the ROB to be sent. This is especially costly during burst accesses. The LLC supporting 16 outstanding requests and having a 10 cycles latency, 16 memory accesses should be pipelined to take a total of $10+16=26$ cycles. Without pipelining, they will now take $10*16=160$ cycles.

### v4

@@ 4.1. Strong Microarchitectural Isolation
Memory Isolation Sanctum (costan2016sanctum, ) introduces hardware mechanisms for memory isolation in the context of an untrusted OS. Like RISC-V Physical Memory Protection (PMP) (waterman2015risc, ), Sanctum introduces a mechanism to partition physical memory and prevent unprivileged security domains (OS or enclave) from accessing the private memory regions of another security domain (OS, enclave or SM). At its core, the mechanism uses two 64-bit registers per core (one for trusted memory and one for untrusted/shared memory) as memory bitmaps. Each bit represents a fixed-size physical memory region (typically 32MB) and functions as an access-rights flag for the currently-running security domain. Citadel borrows this mechanism for its simplicity and enforces these checks on every memory access, including speculative accesses and on accesses performed during page walks. Memory isolation is illustrated in Figure 2, where the coloring highlights that only currently running security domains can access corresponding DRAM memory regions.
Microarchitectural Isolation Following MI6 (bourgeat2019mi6, ), we enforce strong microarchitectural isolation through partitioning, flushing and by disabling speculation in the SM. First, we partition all spatially-shared microarchitecture including the LLC, the LLC pipeline and Miss Status Handling Registers (MSHRs) and other resources as illustrated in Figure 2. In particular, the LLC is statically partitioned into 64 regions to match the 64 DRAM memory regions – we will improve on this scheme in Section 6. We pad the timing of DRAM requests to a constant, which eliminates DRAM controller side channels. When context switching between security domains, we leverage interrupt logic to interpose the SM and enforce the flushing of temporally-shared microarchitecture such as all cores’ pipeline state, L1 caches, TLBs, Translation Caches and MSHRs (see Figure 3). The processor does not support hyper-threading and therefore is not vulnerable to related side channels such as port contention. The processor does not include a prefetcher so we are not susceptible to Spectre variants exploiting them. Finally, Citadel is not susceptible to Meltdown-style attacks, as the processor does not speculate across privilege levels (switching to a different privilege mode flushes the re-order buffer, making sure it never contains two instructions from different privilege levels). The exhaustive list of side channels and defense mechanisms covered by MI6 can be found in (bourgeat2019mi6, ).
Disabling Speculation in the SM The SM executes in machine mode (see Figure 3) which gives it access to all physical memory, making it vulnerable to Spectre-like attacks. Following (bourgeat2019mi6, ), we disable speculation in the SM to prevent these attacks.
@@ 4.2. Enabling Shared Memory
Figure 4. The Dual Page Table Mechanism + Restricting Speculative Access to Shared Memory
Sanctum, a non-speculative design, proposed a dual-page-table mechanism (see Figure 4) to differentiate between trusted and untrusted memory based on virtual addresses and enable shared memory. We adopt this dual-page-table mechanism in Citadel. Two new registers (equivalent to base and bounds) are used to define the trusted virtual memory range (i.e., the enclave’s virtual memory). The SM enforces the invariant that the physical memory mapped as trusted is owned by the corresponding enclave. When a memory access to a virtual address is performed by a program, one of the two page tables needs to be picked to translate the virtual address. If the virtual address falls within the trusted range, the translation is performed using the enclave page table. Another important invariant enforced by the SM is that the enclave page table is always stored within enclave private memory. If the virtual address falls outside the trusted range, it corresponds to an untrusted or shared memory access and the translation is performed using the OS page table. The OS page table is untrusted and stored in OS memory. Note that for every physical address used during address translation, additional bound checks are performed to ensure they belong to the same security domain (enclave or OS memory) as the translated address. This mechanism is especially handy to eliminate speculative transmitters, as there is no need to perform any speculative address translation (which could leak information) to know whether a memory access is trusted or not.
4.3. Attacks and Reusing Existing Mechanisms
Shared memory implies that security domains now share some microarchitectural state, specifically a shared LLC partition. This effectively breaks Strong Microarchitectural Isolation and opens the door to LLC side channels and Spectre-like attacks. Anecdotally, we have successfully mounted such an attack against an enclave on our insecure baseline running on an FPGA. We were not able to successfully mount the attack on our final secure design. To thwart these attacks, we need to achieve SNI (see Section 3) i.e., prevent speculative accesses to shared memory from leaking enclave secrets, for instance, by preventing speculative accesses to shared memory altogether (see Figure 4). Nevertheless, building such a mechanism is not trivial. For instance, existing mechanisms that prevent some speculative memory accesses (MMIO accesses or non-cacheable memory (schwarz2020context, )) cannot be repurposed. Indeed, they rely on physical addresses or page table entries to identify memory regions, implying that they allow some speculative memory accesses to page tables, already leaking part of the address. Instead, we propose to design new mechanisms for speculation-safe shared memory to balance security and performance.
4.4. SpecSafe: Speculatively Safe Shared Memory
@@
Figure 5. Changes to the Memory Execution Pipeline to enable SpecSafe. New elements are shown in blue.
@@
Modifying the Memory Execution Pipeline Upon entry into the MemExePipeline, checks are performed on the memory instruction to decide if the memory operation is safe to translate and then issue, or not. It is important to perform these checks very early as the next steps in the MemExePipeline touch shared microarchitecture (through address translation and cache hierarchy for instance) potentially transmitting information to an attacker. A memory access is deemed safe if it accesses trusted memory OR if it is non-speculative. Because the enclave trusted memory range is defined in virtual memory, this check can be performed as soon as operands (i.e., virtual addresses) are available. If an instruction tries to access untrusted or shared memory, we check that it is non-speculative by conservatively checking if it has reached the head of the ROB. If the instruction is safe, it is dispatched to the next stage of the MemExePipeline. If it is not safe, the instruction is squashed and a signal is sent to the MemRS indicating that it should re-dispatch it once it reaches the head of the ROB.
@@
Modifying the Memory Reservation Station Our mechanism requires MemRS being able to re-dispatch some instructions, once they become non-speculative – untrusted memory loads will indeed typically need to get dispatched twice.
@@
Disabling the Mechanism for the OS Blocking speculation on memory access can have a significant performance impact. The SM makes sure that non-enclave programs (typically the OS and standard user mode programs) are still allowed to speculate when accessing their own memory by disabling the trusted memory mechanism. Note that untrusted code is still barred form (speculatively) accessing enclave or SM memory, thanks to memory isolation (see Section 4.1).
4.5. Burst Mode: Pipelined Shared Memory Accesses
Figure 6. Changes to the Fetch Pipeline to enable Burst Mode. New elements are shown in blue.
@@ 4.5. Burst Mode: Pipelined Shared Memory Accesses
We observe that SpecSafe performs poorly on code that performs back-to-back accesses to shared memory. This is mainly because shared memory accesses cannot be pipelined, as each instruction needs to wait to reach the head of the ROB to be sent. This serialization problem is exacerbated as we cannot use the L1-D cache for security reasons (see Section 6), increasing latency of responses. In some cases, this behavior is even exacerbated by the compiler, trying to optimize the code by placing all the load instructions to shared memory, one after the other. This is especially costly during burst accesses. The LLC supporting 16 outstanding requests and with a 10 cycle latency, means that 16 memory accesses should be pipelined to take a total of $10+16=26$ cycles. Without pipelining, they will now take $10\times 16=160$ cycles.